OpenStudy (anonymous):

The midpoint of is E(–1, 0). One endpoint is C(5, 2). What are the coordinates of the other endpoint?

OpenStudy (anonymous):

pg 51

OpenStudy (anonymous):

since the midpoint is just the average of the coordinates of the endpoints\[-1=\frac{x+5}{2}\]\[-2=x+5\]\[x=-7\] the y-coordinate of the endpoint is\[0=\frac{y+2}{2}\]\[0=y+2\]\[y=-2\]the endpoint is (-7,-2)

OpenStudy (anonymous):

THnak you!!!!

OpenStudy (anonymous):

ur welcome :})